OpenStudy (anonymous):

The perimeter of a triangle is 42 centimeters. The longest side is 2 centimeters less than the sum of the other two sides. Twice the shortest side is 4 centimeters less than the longest side. Find the lengths of each side of triangle please

OpenStudy (anonymous):

Let x denote the length of the longest side and y and z the length of the other 2 sides. Now y+z-2 = x so perimeter = x+y+z = 42 that is x+2+x=42 that is x = 20 and 2y = x-4 that is 2y = 16 that is y = 8 that is z = 14